# Monocular pose estimation of articulated surgical instruments in open surgery

### v1

@@ Abstract
This work presents a novel approach to monocular 6D pose estimation of surgical instruments in open surgery, addressing challenges such as object articulations, symmetries, occlusions, and lack of annotated real-world data. The method leverages synthetic data generation and domain adaptation techniques to overcome these obstacles. The proposed approach consists of three main components: $(1)$ synthetic data generation using 3D modeling of surgical tools with articulation rigging and physically-based rendering; $(2)$ a tailored pose estimation framework combining object detection with pose estimation and a hybrid geometric fusion strategy; and $(3)$ a training strategy that utilizes both synthetic and real unannotated data, employing domain adaptation on real video data using automatically generated pseudo-labels. Evaluations conducted on videos of open surgery demonstrate the good performance and real-world applicability of the proposed method, highlighting its potential for integration into medical augmented reality and robotic systems. The approach eliminates the need for extensive manual annotation of real surgical data.
@@ 1 Introduction
Object pose estimation is a fundamental problem in computer vision that aims to determine the position and orientation of an object in 3D space relative to a camera. It has a wide range of applications, including robotic manipulation Tremblay et al. [57], autonomous navigation [41], and augmented reality [50]. In the medical domain, accurate pose estimation of surgical instruments plays a crucial role in computer-assisted surgery, enabling advanced features such as surgical navigation, skill assessment, and robotic assistance [13, 7, 17].
Traditional approaches to object pose estimation often rely on template matching [58, 45] or key-point matching [31]. However, with the advent of deep learning, significant progress has been made in developing more robust and accurate pose estimation techniques using various input modalities, including RGB images [62, 47], depth maps, and point clouds [59].
@@
Existing approaches to object pose estimation can be broadly categorized into two groups: marker-based [39, 1] and markerless methods. Marker-based methods rely on attaching physical markers, such as color-coded fiducials [61] or passive reflective markers [18], to the instruments. These markers are then detected and tracked in the image space to estimate the pose. While marker-based methods can provide accurate pose estimates, they require modifying the instruments and may interfere with the surgical workflow [10].
@@
On the other hand, markerless methods aim to estimate the pose directly from the visual appearance of the instruments in the image. These methods typically involve training deep neural networks on large datasets of annotated images. The networks learn to extract relevant features and estimate the 6D pose of the instruments using various techniques, such as landmark detection [46] or dense correspondence estimation with Perspective-n-Point (PnP) algorithm [51], or direct pose regression [60]. Markerless methods offer a more practical and non-intrusive solution for surgical instrument pose estimation, as they do not require any modifications to the instruments.
@@
While markerless pose estimation methods have shown promising results, they face several challenges that hinder their widespread adoption in surgical settings. One of the primary challenges is the scarcity of annotated real-world data [3, 9]. Collecting and annotating large datasets of images with accurate 6D pose labels is time-consuming, expensive, and often impractical [25, 57]. This challenge is particularly relevant in the medical domain, where privacy concerns and the difficulty of recording during surgeries due to camera placement restrictions (as some areas must remain sterile) further complicate the data collection process.
@@
To address the data scarcity problem, researchers have explored the use of synthetic data generation techniques [56, 57]. By creating realistic virtual models of surgical instruments and rendering them in simulated environments, large amounts of annotated training data can be generated efficiently. However, models trained solely on synthetic data often suffer from the domain gap problem, where the learned features do not generalize well to real-world scenarios. To bridge this gap, domain adaptation techniques, such as domain randomization [57] and adversarial training [4], have been proposed. These techniques aim to make the models more robust to the differences between synthetic and real data, improving their performance on real-world images.
@@
Despite the significant progress in deep learning-based pose estimation methods, their application to surgical instruments poses additional challenges due to the unique characteristics of the surgical environment and the instruments themselves [42, 15]. Surgical instruments often have metallic and reflective surfaces, which can cause severe specular reflections and make it difficult to extract reliable visual features. Moreover, articulated objects, such as surgical instruments with movable parts, introduces additional complexities. Articulated objects have multiple degrees of freedom, allowing their parts to move independently, which significantly increases the difficulty of accurately estimating their pose.
@@
In this study, we tackle the multiple challenges raised by of open surgery pose detection. Our key contributions include:
@@
    Utilization of solely synthetic data to develop methods for both object detection and tool pose estimation. The pose estimation incorporates the standard six degrees of freedom as well as ool articulation (i.e., measuring how open the tools are).
@@
    Introduction of a sim2real domain adaptation technique that utilizes the temporal-spatial characteristics of video data to refine both object detection and pose estimation models, with each model helping to improve the other.
It is important to highlight that our approach eliminates the need for manual data labeling. New tools can be added to the dataset using only their 3D models (e.g., STL files) without manual labeling. Any labeling done in this study serves solely to validate our algorithms. Addressing these challenges is crucial for developing accurate and reliable pose estimation methods for surgical instruments, which can ultimately lead to improved surgical outcomes and patient care.
@@ 1.1 Object Pose Estimation Methods
Object pose estimation methods have seen significant advances in recent years, particularly with the rise of deep learning. In this section, we discuss state-of-the-art methods for both general object pose estimation and those specifically designed for surgical instruments. Single-stage methods directly regress the 6D pose of an object from the input image. [62] predicts the 3D translation and rotation of an object in a single forward pass. Having separate branches for object classification and pose estimation. The pose estimation branch regresses the 3D translation and rotation in a quaternion representation. [12], extends the popular object detection framework Mask R-CNN [22] to jointly perform object detection and pose estimation. It introduces a pose estimation branch that operates on the region of interest(ROI) features extracted by the ROI pooling layer. It combines object detection and pose estimation in a single network, allowing pose estimation of multiple objects in an image. [64] is a single-stage method that estimates 2D-3D dense correspondence maps on the full image and estimates the pose via PnP-RANSAC. Then, further refine the initial pose estimate using a custom deep learning model.
Two-stage methods decouple object detection and pose estimation by first detecting the object of interest and then estimating its pose using a separate network. [47] employs a segmentation-based approach to identify the 2D centers of the objects of interest. The segmentation network does not use a hourglass-shaped architecture, making it more efficient. It then predicts the 3D bounding box of the object and regresses the 6D pose from the cropped object image. The method employs a coarse-to-fine approach, where the initial pose estimate is refined using a more precise network.
[32] disentangles the pose estimation process into predicting the object’s rotation and translation separately. This disentanglement reportedly leads to more accurate and robust pose estimation compared to previous methods.
[44] employs a fully convolutional network to predict dense 2D-3D correspondences and utilizes a GAN to enhance the stability of these dense correspondences.
[46] presents a voting-based approach to manage occluded or truncated keypoints. Each pixel votes for the object’s keypoints, and the pose is determined by solving a least-squares problem. The network generates pixel-wise heatmaps for each keypoint and employs a differentiable RANSAC layer to estimate the pose from the keypoint correspondences.
Several state-of-the-art deep learning architectures have been proposed to further improve the accuracy and robustness of object pose estimation.
[51] focuses on surface encoding by assigning a unique code to each vertex on a 3D object model, starting with coarse fragments and refining to finer details. The method includes assigning these unique codes to 3D vertices, predicting their correspondence to image pixels, and estimating the object’s pose using a PnP-RANSAC-like approach.
[60] employs a direct regression approach to efficiently and differentiably predict the 6D pose of an object in an end-to-end trainable manner. By utilizing intermediate geometric representations, the method guides and enhances the accuracy of pose prediction. It achieves competitive performance compared to state-of-the-art methods on standard benchmarks.
While these methods have shown impressive results on general object pose estimation benchmarks, they often struggle when applied directly to surgical instruments due to the unique challenges posed by the surgical environment.
1.2 Pose Estimation Methods for Surgical Instruments
Several deep learning-based methods have been proposed specifically for surgical instrument pose estimation, addressing the unique challenges posed by the surgical environment. These methods can be further categorized based on the type of surgery they focus on, such as laparoscopic or open surgery.
Methods Focusing on Laparoscopic Instruments. [21] present a method for detection, segmentation, and 3D pose estimation of laparoscopic instruments using CNN’s and algebraic geometry. They employ a two-stage approach, where the instrument is first detected and segmented using a CNN-based segmentation network. The segmented instrument is then processed by a separate CNN to predict its 3D pose using an algebraic geometry-based method. The authors demonstrate the effectiveness of their approach on a dataset of laparoscopic images, achieving high accuracy in instrument detection, segmentation, and pose estimation.
Methods addressing challenges in open surgery. [23] introduce a markerless hand-tool tracking pipeline for open surgery, focusing on larger, non-articulated tools. They create a synthetic data generation pipeline using the MANO hand model [49] and GraspIt! [40] simulator to produce realistic hand-tool interactions. The authors also capture real-world data with ground truth annotations in a mock operating room. They evaluate three baseline models ([46], HandObjectNet, and a combined model) trained on the synthetic dataset and refined on the real dataset. The best-performing model, HandObjectNet, achieves an average 3D vertex error of 16.7 mm on the synthetic test set and 13.8 mm on the real test set after refinement. This work demonstrates the potential of using synthetic data for pretraining and real-world data for fine-tuning pose estimation models in open surgery.
1.3 Overview of Synthetic Data Generation for Object Pose Estimation
Synthetic data generation has gained significant attention in the object pose estimation community due to several reasons. First, annotating real-world data for object pose estimation is labor-intensive and error-prone, often requiring specialized equipment or manual labeling by experts. In contrast, synthetic data can be automatically generated with precise annotations, saving time and resources. Second, rendering synthetic data is fast and cheap compared to capturing real-world data, allowing researchers to create large-scale datasets with diverse object appearances, poses, backgrounds, and lighting conditions. This diversity can improve the robustness and generalization of trained object pose estimation models.
One of the most common approaches for generating synthetic data for object pose estimation is through 3D modeling and rendering. This involves creating 3D models of objects using computer graphics software and rendering them from different viewpoints to generate 2D images with corresponding pose annotations. The 3D models can be created manually using 3D modeling software or scanned from real objects using techniques like photogrammetry or depth sensing.
To create realistic synthetic images, physically-based rendering (PBR) techniques are often employed. PBR accurately simulates the flow of light energy in the scene by ray tracing, naturally accounting for complex illumination effects such as scattering, refraction, and reflection [26]. The resulting rendered images look realistic and are often difficult to differentiate from real photographs.
In the context of object pose estimation, 3D models of objects are placed in virtual scenes and rendered from various viewpoints. The precise 3D poses of the objects in these rendered images can be automatically recorded, providing accurate ground-truth annotations for training and evaluating object pose estimation models. By varying the object poses, backgrounds, and lighting conditions in the virtual scenes, researchers can generate diverse synthetic datasets that cover a wide range of scenarios.
The effectiveness of using synthetic data generated through 3D modeling and PBR for object pose estimation has been demonstrated in various studies. In the BOP challenge 2020 on 6D object localization [26], methods trained on PBR images achieved significantly higher accuracy scores compared to those trained on ”render & paste” images. While incorporating real training images further improved the scores, competitive results were obtained using only PBR images. Notably, the increased photorealism of PBR images led to clear improvements even for methods that apply strong data augmentation, such as CosyPose [30].
1.4 Bridging the Domain Gap Between Synthetic and Real Data
@@
While synthetic data generation has shown great promise in training deep learning models for object pose estimation, there often exists a domain gap between synthetic and real-world data [57]. This domain gap can limit the performance of models trained solely on synthetic data when applied to real-world scenarios. Various techniques have been proposed to address this issue to bridge the domain gap and improve the transferability of models trained on synthetic data to real-world applications.
Domain adaptation techniques:
  1. 1.
    Fine-tuning: One common approach is to fine-tune models that are pre-trained on synthetic data using a smaller dataset of real-world images. This allows the models to adapt to the characteristics of real-world data while leveraging the knowledge learned from synthetic data. Fine-tuning can be performed on the entire model or only on specific layers, depending on the similarity between the synthetic and real-world domains [55, 56].
  2. 2.
    Unsupervised domain adaptation: Unsupervised domain adaptation techniques aim to align the feature distributions of synthetic and real-world data without requiring labeled real-world examples. These techniques often involve adversarial training, where a discriminator network is trained to distinguish between features from the synthetic and real-world domains, while the pose estimation network is trained to fool the discriminator. By minimizing the domain discrepancy, unsupervised domain adaptation can improve the performance of object pose estimation models on real-world data [20].
@@
Examples of bridging the domain gap: Several works have successfully demonstrated the effectiveness of bridging the domain gap between synthetic and real data for object pose estimation. For instance, the work by [23] on hand-tool interaction in open surgery utilizes a combination of synthetic pretraining and real-world fine-tuning. They first train their models on a large-scale synthetic dataset generated using a physically-based rendering pipeline and then fine-tune the models on a smaller dataset of real-world images captured in a mock operating room. This approach allows them to leverage the diversity and accuracy of synthetic data while adapting the models to the characteristics of real-world surgical scenes.
@@
Another example is the work by Sundermeyer et al. [53], where they propose a self-supervised domain adaptation method for object pose estimation. Their approach relies on a differentiable renderer to generate synthetic views of objects and align them with real-world images using a contrastive loss. By minimizing the domain discrepancy between synthetic and real-world images, their method can improve the performance of object pose estimation models on real-world data without requiring any labeled real-world examples.
@@ 1.5 Data Collection
We collected real-world data from surgical procedures for the purpose of unsupervised sim-to-real domain adaptation and to validate our approach and ensure its effectiveness in practical surgical scenarios. Data were collected from seven different surgeries. Each surgery involved the removal of a vein from the leg as part of the “Coronary Artery Bypass Graft (CABG)” procedure. Our focus is specifically on the suturing part of the procedure, which typically lasts between 15 to 30 minutes. In total, four different surgeons performed the surgeries, with each suturing procedure carried out by a single surgeon. The study was approved by the Rambam Health Care Campus IRB committee.
@@ 1.6 Data Annotation for Evaluation
Since annotating 6D pose is very complex, we instead annotate instance segmentation labels and evaluate on them. Approximately 100 images were selected from each procedure, covering a variety of tool poses and challenging scenarios. These frames were manually annotated.
For each selected frame, we created detailed segmentation masks of the surgical tools of interest (needle-holder and tweezers), as well as background surgical tools and hands. These segmentation masks provide pixel-level annotations of the tools and hands in the surgical scene. The annotation process was carried out using the SAM (Segment Anything) model [28], which generates high-quality object masks from input prompts such as points or boxes. Any masks that did not meet quality standards were manually adjusted. These annotated frames provide essential ground truth data for evaluating the effectiveness of our pose estimation network on actual surgical data.

### v2

@@ Abstract
This work presents a framework for monocular 6D pose estimation of surgical instruments in open surgery, addressing challenges such as object articulations, specularity, occlusions, and synthetic-to-real domain adaptation. The proposed approach consists of three main components: $(1)$ synthetic data generation pipeline that incorporates 3D scanning of surgical tools with articulation rigging and physically-based rendering; $(2)$ a tailored pose estimation framework combining tool detection with pose and articulation estimation; and $(3)$ a training strategy on synthetic and real unannotated video data, employing domain adaptation with automatically generated pseudo-labels. Evaluations conducted on real data of open surgery demonstrate the good performance and real-world applicability of the proposed framework, highlighting its potential for integration into medical augmented reality and robotic systems. The approach eliminates the need for extensive manual annotation of real surgical data.
keywords:
[…]
@@ 1 Introduction
Object pose estimation is a fundamental problem in computer vision that aims to determine the position and orientation of an object in 3D space relative to a camera. It has a wide range of applications, including robotic manipulation [tremblay2018deep], autonomous navigation [mousavian20173d], and augmented reality [su2019deep]. In the medical domain, pose estimation of surgical instruments plays a crucial role in computer-assisted surgery, skill assessment, and robotic assistance [burton2023evaluation, bkheet2023using, doignon2008pose].
@@
Existing approaches to object pose estimation can be broadly categorized into two groups: marker-based [meza2021markerpose, ababsa2004robust] and markerless methods. Marker-based methods rely on attaching physical markers, such as color-coded fiducials [wei1997real] or passive reflective markers [elfring2010assessment], to the instruments. These markers are then detected and tracked in the image space to estimate the pose. While marker-based methods can provide accurate pose estimates, they require modifying the instruments and may interfere with the surgical workflow [bouget2015detecting].
@@
On the other hand, markerless methods aim to estimate the pose directly from the visual appearance of the instruments in the image. These methods typically involve training deep neural networks on large datasets of annotated images. The networks learn to extract relevant features and estimate the 6D pose of the instruments using techniques such as correspondence estimation with Perspective-n-Point (PnP) algorithm [su2022zebrapose, wang2021gdr]. Markerless methods offer a more practical and non-intrusive solution for surgical instrument pose estimation, as they do not require any modifications to the instruments.
@@
While markerless pose estimation methods have shown promising results, they face several challenges that hinder their widespread adoption in surgical settings. One of the primary challenges is the scarcity of annotated real-world data [allan20192017, bouget2017vision]. Collecting and annotating large datasets of images with accurate 6D pose labels is time-consuming, expensive, and often impractical [hodan2018bop, tremblay2018deep]. This challenge is particularly relevant in the medical domain, where privacy concerns and the difficulty of recording during surgeries due to camera placement restrictions (as some areas must remain sterile) further complicate the data collection process.
@@
To address the data scarcity problem, researchers have explored the use of synthetic data generation techniques [tremblay2018training, tremblay2018deep]. By creating realistic virtual models of surgical instruments and rendering them in simulated environments, large amounts of annotated training data can be generated efficiently. However, models trained solely on synthetic data often suffer from the domain gap problem, where the learned features do not generalize well to real-world scenarios. To bridge this gap, domain adaptation techniques, such as domain randomization [tremblay2018deep] and adversarial training [bai2021recent], have been proposed. These techniques aim to make the models more robust to the differences between synthetic and real data, improving their performance on real-world images.
@@
Despite the significant progress in deep learning-based pose estimation methods, their application to surgical instruments poses additional challenges due to the unique characteristics of the surgical environment and the instruments themselves [nema2022surgical, colleoni2020synthetic]. Surgical instruments often have metallic and reflective surfaces, which can cause severe specular reflections and make it difficult to extract reliable visual features. Moreover, articulated objects, such as surgical instruments with movable parts, introduces additional complexities. Articulated objects have additional degrees of freedom due to their moving parts, which significantly increases the difficulty of accurately estimating their pose.
@@
Our key contributions include:
@@
    Pose estimation network designed for articulated surgical tools without predefined geometric assumptions.
@@
    Introduction of a synthetic-to-real domain adaptation technique that leverages the temporal features of video data to enhance both object detection and pose estimation models, enabling mutual improvement between the two.
It is important to emphasize that our approach removes the requirement for manual data labeling. New tools can be incorporated into the dataset using only their 3D models (e.g., STL files), without the need for manual annotation. Any labeling performed in this study is strictly for the purpose of validating our algorithms.
@@ 1.1 Object Pose Estimation Methods
The advent of deep learning has propelled significant advancements in object pose estimation, giving rise to diverse methodologies. These approaches can be broadly categorized into single-stage versus two-stage methods, direct pose regression techniques, and keypoint detection followed by PnP-RANSAC. These categories are not mutually exclusive, and many state-of-the-art methods combine elements from different approaches to achieve better performance.
Single-stage methods [xiang2017posecnn, braun2016pose, he2017mask, zakharov2019dpod], integrate object detection and pose estimation within a unified network, offering a streamlined process. This approach allows for end-to-end training and potentially faster inference times. In contrast, two-stage methods [li2019cdpn, park2019pix2pose, peng2019pvnet, su2022zebrapose, wang2021gdr], decouple the object detection and pose estimation tasks. While this separation may increase computational complexity, it allows for independent optimization of each stage, potentially leading to improved accuracy.
Keypoint detection methods can be further categorized into sparse and dense techniques [peng2019pvnet, wang2021gdr, su2022zebrapose]. Sparse keypoint detection identifies a limited number of significant points on the object within the image. Dense keypoint detection recognizes a more extensive set of points across the object’s surface in the image, potentially providing richer information for pose estimation. For instance, [peng2019pvnet] employs a voting-based method to handle occluded or truncated keypoints, demonstrating the robustness of keypoint-based approaches.
After keypoint identification, pose estimation techniques generally fall into two categories: Classical PnP-RANSAC [peng2019pvnet, su2022zebrapose] and direct pose regression [wang2021gdr]. The classical PnP-RANSAC approach uses the identified keypoints to solve the Perspective-n-Point (PnP) problem, often combined with RANSAC for robust estimation. Direct regression techniques aim to predict pose parameters in an end-to-end fashion.
Recent works have introduced innovative approaches. [su2022zebrapose] focuses on dense surface encoding by assigning unique codes to 3D object model vertices and then uses a PnP-RANSAC variant to determine the pose. [wang2021gdr] utilizes dense intermediate geometric representations and employs a direct regression method to estimate the pose.
While these methods have shown impressive results on general object pose estimation benchmarks, they often face challenges when applied to surgical instruments due to the unique characteristics of the surgical environment. In response to these challenges, researchers have developed specialized methods for surgical instrument pose estimation. For laparoscopic instruments, [hasan2021detection] proposed a two-stage approach. Their method employs a CNN-based segmentation network for instrument detection and segmentation, followed by a separate CNN for 3D pose estimation using algebraic geometry. This approach demonstrated high accuracy on a dataset of laparoscopic images. For open surgery scenarios, [hein2021towards] introduced a markerless hand-tool tracking pipeline, focusing on larger, non-articulated tools. Their approach involved creating synthetic data using the MANO hand model [MANO:SIGGRAPHASIA:2017] and GraspIt! simulator [miller2004graspit], capturing real-world data in a mock operating room for fine-tuning, and evaluating multiple baseline models. In a subsequent multi-view approach [hein2023next] introduced a multi-camera capture setup (including both static and head-mounted cameras) and a multi-view RGB-D dataset of ex-vivo spine surgeries, evaluating single- and multi-view pose estimation methods for surgical instrument tracking. Their results demonstrated sub-millimeter accuracy for certain instruments under optimal conditions, suggesting that markerless, multi-view tracking can be a feasible alternative to existing marker-based systems.
Synthetic data generation has gained significant attention in the object pose estimation community. The common approach involves 3D modeling and rendering, using physically-based rendering (PBR) techniques to create realistic images with accurate pose annotations. The effectiveness of this approach was demonstrated in the BOP challenge 2020, where methods trained on PBR images significantly outperformed those trained on simpler ”render & paste” images.
@@
While synthetic data generation has shown great promise in training deep learning models for object pose estimation, there often exists a domain gap between synthetic and real-world data [tremblay2018deep]. Various domain adapatation techniques have been proposed to bridge this domain gap and improve the transferability of models trained on synthetic data to real-world applications.
@@
Several works have successfully demonstrated the effectiveness of bridging the domain gap between synthetic and real data for object pose estimation. For instance, the work by [hein2021towards] on hand-tool interaction in open surgery utilizes a combination of synthetic pretraining and real-world fine-tuning. They first train their models on a large-scale synthetic dataset generated using a physically-based rendering pipeline and then fine-tune the models on a smaller dataset of real-world images captured in a mock operating room. This approach allowed them to leverage the diversity and accuracy of synthetic data while adapting the models to the characteristics of real-world surgical scenes.
Similarly, [souipas2023simps] achieve markerless tool segmentation and 3D localization without requiring prior knowledge of tool structure. Although they focused on articulated surgical tools, they did not explicitly handle the articulations and conducted their experiments on mockup operations.
@@
Another example is the work by [sundermeyer2020augmented], where they propose a self-supervised domain adaptation method for object pose estimation. Their approach relies on a differentiable renderer to generate synthetic views of objects and align them with real-world images using a contrastive loss. By minimizing the domain discrepancy between synthetic and real-world images, their method can improve the performance of object pose estimation models on real-world data without requiring any labeled real-world examples.
2 Data Collection and Annotation
For this work, we collected real-world data from surgical procedures for the purpose of unsupervised sim-to-real domain adaptation and to validate our approach and ensure its effectiveness in practical surgical scenarios.
@@ 2.1 Data Collection
We collected data from ten different surgeries. Each surgery involved the removal of a vein from the leg as part of the ”Coronary Artery Bypass Graft (CABG)” procedure. Our focus is specifically on the suturing part of the procedure, which typically lasts between 15 to 30 minutes. In total, five different surgeons performed the surgeries, with each suturing procedure carried out by a single surgeon. The study was approved by the Rambam Health Care Campus IRB committee.
@@ 2.2 Data Annotation for Evaluation
To ensure unbiased evaluation, three surgical videos were reserved exclusively for domain adaptation and the other seven for testing. Approximately 100 images were selected from each testing procedure for evaluation, covering a variety of scenarios.
Since annotating 6D pose is very complex, we instead annotate instance segmentation labels and evaluate on them. For each selected frame, we created detailed segmentation masks of the surgical tools of interest (needle-holder and tweezers), as well as background surgical tools and hands. These segmentation masks provide pixel-level annotations of the tools and hands in the surgical scene. The annotation process was carried out using the SAM (Segment Anything) model [kirillov2023segany], which generates high-quality object masks from input prompts such as points or boxes. Any masks that did not meet quality standards were manually adjusted. These annotated frames provide essential ground truth data for evaluating the effectiveness of our pose estimation network on actual surgical data.